Normalize the following untyped applied lambda calculus expressions as much as possible using the call-by-value rules, showing all steps CLEARLY, please. TYPED answer is preferred.

a) (λx. x * x) 1

b) (λx. x + 4) ((λy. y + 5) 3)

c) (λf g x. g (f x)) (λa. a * a) (λb. b + 1 + 2) 3

d) (λf x. f (f (f x))) (λb. if b then false else true) true
Answer #1

assume λx.R .....where R is lambda expression.

a) (λx. x * x) 1

Solution: uderline show the lambda expression

put 1 in the lambda expression

x*x = 1*1 = 1 Ans

b) (λx. x + 4) ((λy. y + 5) 3)

solution: step1:  

((λy. y + 5) 3) ----> y + 5) 3) lambda expression

f(x) value is 3 in place of x....

In given expression write such as f(y) = y+3, where y = 3

so value of expression is 3+5 =8

step : now comlete expression reduced to

(λx. x + 4)8

lambda expression is (x + 4)8 so we put value 8 in place of x

step

8+4 = 12 Ans

c) Solution:

(λf g x. g (f x)) (λa. a * a) (λb. b + 1 + 2) 3

Step   (λb. b + 1 + 2) 3 solve first

step

(b + 1 + 2) 3 = 3+1+2 =6

step solve   (λa. a * a)6

(a * a)6 = 6*6 =36

step solve (λf g x. g (f x))36